OpenStudy (anonymous):

Solve the following equation. Be sure to check your work. (28b -7) -9(3b+1)= -3

OpenStudy (anonymous):

like terms all go first

OpenStudy (anonymous):

This is what I got so far

OpenStudy (anonymous):

I mean all ways so - 28b -3b

OpenStudy (anonymous):

28b -7 -27b -9 = -3

OpenStudy (anonymous):

(28b-7)-9(3b+1)= -3 28b-7-27b-9=-3 b-16=-3 b=13

OpenStudy (anonymous):

25b

OpenStudy (anonymous):

yea I agree 13

OpenStudy (anonymous):

So 13 is the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

I think the answer is negative

OpenStudy (anonymous):

you can plug 13 back into the equation for b to double check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no it not it a postive